OpenStudy (anonymous):

what is the derivative of ln(x^2) ??? and why???

OpenStudy (anonymous):

2/x

OpenStudy (anonymous):

what happens to ln?

OpenStudy (amistre64):

ln(x^2) = 2ln(x) -> 2/x

OpenStudy (amistre64):

it derives

OpenStudy (anonymous):

ahhh i see

OpenStudy (anonymous):

then how do u find the derivative of (lnx)^4??

OpenStudy (anonymous):

do u put 4 in front of lnx ??

OpenStudy (amistre64):

\[\lim_{h\to o}\frac{ln(x+h)-ln(x)}{h}\] \[\lim_{h\to o}\frac{ln(\frac{x+h}{x})}{h}\] \[\lim_{h\to o}\frac{1}{h}ln(1+\frac{h}{x})\] \[\lim_{h\to o}\ ln(1+\frac{h}{x})^{1/h}\] you never really do the first principles afterwards do ya ...

OpenStudy (amistre64):

when you learn the power rule x^4; you do, 4x^3 the power rule should be taught as functions instead ....

OpenStudy (anonymous):

but im looking for derivative not limits

OpenStudy (amistre64):

(f(x))^4 = 4(f(x))^3 * f'(x)

OpenStudy (amistre64):

derivatives ARE limits .....

OpenStudy (amistre64):

the rules for derivatives are the result of taking their limits ....

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

what happens to ln in that problem?

OpenStudy (amistre64):

it pops out as a chain rule

OpenStudy (amistre64):

(ln(x))^4 -> 4(ln(x))^3 * ln'(x)

OpenStudy (anonymous):

dont u divide the problem by 3 cuz of the exponent?

OpenStudy (amistre64):

i dont see why you would have to make up new rules that have no bearing on the problem

OpenStudy (amistre64):

when we integrate we divide by the exponent; when we take the derivative we mutliply by the exponent

OpenStudy (anonymous):

yea thats what im saying didnt we just integrate?

OpenStudy (amistre64):

you asked: "then how do u find the derivative of (lnx)^4?? "

OpenStudy (amistre64):

so tell me, are we integrating or taking a derivative?

OpenStudy (anonymous):

ohh oops

OpenStudy (anonymous):

i got the answer thx